# Zebrafix: Mitigating Memory-Centric Side-Channel Leakage via Interleaving

### v1

@@ 7.2 Comparison of Zebrafix to Silent Store Mitigations
The Zebrafix implementation protects against leakage from silent stores without requiring further adaptions as long as the silencing of writes cannot be decomposed into smaller chunks. We therefore compare Zebrafix to the silent store mitigation cio [19] that has also been implemented for x86-64. The Zebrafix implementation addresses the general root cause of memory-centric leakages. In line with that, the proof-of-concept implementation does not include a tracking for leaking or secret-dependent instructions, but rather instruments all memory writes (see more in Section 9.1). The silent store-only version of cio only covers a subset of all writes that are instrumented in Zebrafix. Zebrafix does, in contrast to cio, not have an automated verification mechanism but rather introduces manual verification of the non-leakiness of memory writes that are introduced by the register allocation.
@@
Out of the targets evaluated in cio, we compare against $\mathtt{Ed25519}$ and $\mathtt{ChaCha20}$ which were also shown to be (partially) vulnerable in [7]. For $\mathtt{ChaCha20}$, we refer to the encryption; the decryption results are similar. The remaining targets evaluated in cio are not included in this case study because they contain instructions that are not supported in our proof-of-concept due to the way the interleaving is implemented based on vector register writes. Yet, the analyzed primitives are implemented to be executed in vector registers so that the leaking points are (key) initialization and memory cleanup. It is possible to manually safeguard this limited set of code.
@@
The performance results for cio are based on results reported in their paper. In line with their work, we use $\mathtt{libsodium}$ reference implementations and protect stores of up to 8 byte of data. Protecting a binary against silent store leakage with cio introduces an overhead of $1.85$$\times$ for $\mathtt{ChaCha20}$ and an overhead of $3.76$$\times$ for $\mathtt{Ed25519}$. In contrast to that, the ciphertext side-channel protection from Zebrafix introduces an overhead factor of $2.1$$\times$ for $\mathtt{ChaCha20}$ and a factor of $1.3$$\times$ for $\mathtt{Ed25519}$.
@@
In line with [7], we found certain possible leakages in the $\mathtt{ChaCha20}$ target. However, those leakages are not originated in multiple systematic collisions as in $\mathtt{cmov}$ or $\mathtt{cswap}$ functionality but rather single collisions with a large set of possible values that are written during each execution. According to current knowledge, we do not deem those writes as exploitable which is also in line with cio being able to prune a large set of memory writes from their instrumentation.
@@
8 Preventing Leakages from DMP
With suitable instrumentation in place, all memory-centric leakages that arise from writing repetitive data to memory (see Section 2) can be avoided. However, there also are tight relations between ciphertext side-channels and read-based leakages like data memory-dependent prefetching (DMP). In this section, we analyze interconnections between ciphertext side-channel and DMP attacks, and outline to what extent interleaving may be adapted for countering DMP leakage. Our threat model assumption for these considerations is the same as in Section 7.
8.1 Ciphertext Side-Channel Mitigation vs. DMP
The key idea to preventing DMP leakage is breaking the observable relation between data and the corresponding DMP activation. Ciphertext side-channel mitigations that add masking to all secret plaintexts before writing them to memory are thus effective, if they randomize the underlying data. However, for providing a sufficient security level, existing masking approaches like Cipherfix incur high overheads due to the expensive random number generation [42]. The 16-byte block interleaving in Zebrafix that promises lower performance overhead can still not be used directly: Pointers are typically aligned to 8-byte boundaries, so they are unaffected by the additional counter in every second 8-byte chunk. If the DMP still triggers for the new memory layout, it can directly dereference these pointers, ignoring the counters altogether.
8.2 Extending Interleaving to DMP: Approach
If interleaving is to be used for DMP prevention, a smaller granularity is needed to break up DMP activation patterns, effectively reducing the block size from 16 bytes to 8 bytes. Instead of dividing data into 8-byte chunks and interleaving them with 8-byte counters, the data and counter chunk sizes need to be shrunk to implementation-dependent smaller chunks, such as 4 bytes each. For pure DMP mitigation, the counters do not need to be incremented and can be set to a fixed value that prevents DMP activation. If the mitigation is meant also protect against other classes of memory-centric side-channel leakage, the counters can be incremented as before while ensuring that no values occur that again trigger the DMP. An illustration of the new memory layout can be found in Figure 8.
In this mitigation description, we assume that only canonical user-level pointer candidates, i.e., virtual addresses that may trigger prefetching, leak secret information via getting dereferenced. This is in line with current findings [39], that exclude leakage of non-canonical or kernel space addresses. With interleaving-based methods, mitigations may targets the root cause of the leakage, i.e., the DMP activation on a data chunk. Thereby, the aim is to also prevent subsequent exploitable leakages stemming from side-channels based on address translation or TLB lookups [39]. In addition, we assume the worst case leakage scenario, i.e., that the microarchitecture is already in a state that would immediately leak information when a pointer candidate is selected for prefetching.
In the analysis of the DMP in the newer Intel Raptor Lake processors, Wang et al. state that the DMP the dereference targets must reside within the same 256 KB region as the load address [39]. If an interleaving-based mitigation fills the necessary bytes of the potential pointer with random values, there should be no direct leakage of secrets in each of the scanned 8-byte-aligned pointer chunks. Note that we base the explanation of extensibility on the attack vectors presented in [39], combined with theoretical reasoning about address translation properties. Concrete implementations of microarchitectural optimizations may vary regarding specific processor or architectural features.
8.3 Implementing Interleaving for DMP: Sketch
We now outline the necessary changes to the interleaving approach, as proposed in Section 3.2, to also prevent DMP leakage. As mentioned before, the block size of the interleaved structs has to be reduced from 16 byte to the new block size of 8 byte. By that, the supported data size is reduced to 4 bytes, which causes significant problems: 8-byte integers, which are mandated by the architecture and widely used in cryptographic libraries, need to be split into two 4-byte parts. Such split integers can no longer be read through a standard memory load, but need to be emulated. With that, a full pointer tracking becomes necessary, in order to avoid functional crashes of the binary. Additionally, the changes affect a primitive type that is assumed to exist throughout the compiler framework, so tweaks need to be applied at all layers. Due to the significant engineering effort, we leave the implementation of the exploration of this interleaving adaption for future work.
In addition to supporting another block size, the former freshness value has to be replaced by a combination of counter and DMP invalidation. The counter cannot be arbitrarily chosen, as it has to differ in the necessary address bits, putting the data outside the range that is used for determining pointer candidates. Memory writes work similar to interleaving, just with the reduced counter and data size. An implementation needs to be careful with endianness, as to ensure that the counter ends up transforming the necessary bytes of a potential pointer. Address calculations have to be adjusted similarly than memory reads so that the data is loaded correctly for different types of source and result elements of the calculation.
Observation 3 Learning from other memory-centric side-channel mitigations through interleaving can inspire further mitigations against data-at-rest leakage.
9 Discussion
From our observations, we conclude that addressing the root cause of memory-centric side-channel leakages can be based on compiler middle end memory layout adjustments to ensure freshness, as well as improvements by learning from other mitigations in similar areas. In the following, we discuss exploitability metrics for memory-centric side-channel leakages and lessons learned from the implementation of ciphertext side-channel mitigations.
9.1 Leakage Exploitability Estimation
In order to make the mitigation introduced by Zebrafix more efficient, instrumentation could be performed only on potentially vulnerable instructions instead of every memory write. For that, a precise analysis of leaking instructions with a suitable exploitability metric is needed. In the case of ciphertext side-channel leakage, CipherH found some potentially leaking instructions with the help of symbolic execution. Yet, as described in Section 7.1, the proposed approach does not cover all potential leakages but only a subset that would leak in the silent store leakage model. Therefore, in the case of ciphertext side-channel leakage, it would be more secure to overestimate the leaking instructions and take all secret-dependent writes into consideration.
Silent store leakage can be modeled with approaches that check two consecutive memory writes without taking other write accesses to the same address into account. The modeling has also been adapted to leakage models that do not only take code-based symbolic constraints [15] into account but rather a more general modeling of microarchitectural optimizations via domain-specific language-based leakage clauses [7].
Regarding leakage through DMP, it is important to note that all secret-dependent memory writes can potentially harm the security guarantees of even constant-time implementations, as the side-channel can span multiple layers of indirection via pointer dereferencing in the attacker-controllable data. Consequently, if all the leakages discussed in Section 2 are to be mitigated by Zebrafix, the maximum possible improvement is the inclusion of secret tracking, which yields the greatest possible efficiency gain while preserving security guarantees.
9.2 Lessons Learned for Future Mitigations
@@
As described in Section 4, instrumenting on compiler middle end level has multiple advantages and disadvantages. One of the main disadvantages of any static instrumentation method is that language features, such as determining the size of a buffer by subtracting pointers, cannot be instrumented semantically correctly, and must be manually replaced in the source code.
Another common issue during our implementation was the evaluation of $\mathtt{sizeof}$, which already happens in the compiler frond end. This occurred in cases where $\mathtt{sizeof}$ was used to determine the input size of operations like $\mathtt{memcpy}$ or $\mathtt{memset}$, which work on arrays. In our proof-of-concept implementation of Zebrafix, we work around that problem by retrieving the correct byte count from the LLVM $\mathtt{DataLayout}$ object and combining that with type information from the SVF pass.
We support memory writes up to 8 bytes in size, which is the integer size on x86-64. For larger writes (e.g., vector instructions), the data would have to be split into chunks and written within multiple interleaved chunks. Although we have not implemented vector support in Zebrafix, we do not expect any specific performance impact when rewriting vectorized stores.
@@
Most cryptographic libraries are implemented in C, a language that lacks type safety, which poses certain key challenges for memory layout adjustments. Some manual intervention is needed for functions that use “indirect” casts on input data like putting a char array into a function that expects to work with 4-byte integers. Such casts might lead to semantically wrong results when applying operations that depend on the data type’s width like copying data from one buffer into another. The default handling case can only assume 8-byte data chunks in order to avoid out-of-bounds reads or writes, so any function that displays a warning has to be checked and then adjusted in the LLVM SVF pass handling for the correct pointer tracking. The currently implemented solution relies on the fact that function names are available in such cases for easing manual adjustments. With C, there is the choice between manual code adaptation, significant compiler code analysis, or extensive runtime type tracking. Alternatively, one can rely on a language in which such casts do not occur, and it is always known what binary type a variable has.
@@
Another general issue arising in the proposed instrumentation approach stems from global variable usage: Either all code that uses global variables can be instrumented or no code at all. In case a mitigation should be implemented with additional secrecy tracking, the correct functionality cannot be guaranteed for global variable usage. If instrumented as well as non-instrumented parts of the code used global variables, the global variable state might destroy synchronization of values between different functions. Moreover, the general functionality can be compromised as even if all global variable usages were replaced with the instrumented version, there are usages depending on the global variables that cause problems in function parameters or functions like $\mathtt{memcpy}$ or $\mathtt{memset}$ or in $\mathtt{GEP}$ accesses for address calculations. [^5] Even with optimization level $\mathtt{-O2}$ and having all GEP accesses combined in one instruction due to the LLVM InstCombine pass, the iteration steps in arrays can still be wrong if the source element type of a GEP is not interleaved while the pointer of it points to an interleaved (instrumented) global variable. Further analyses would incur large overheads and make the resulting binary less stable; Apart from that, we deem the scenario of needing a non-instrumented version of each function alongside of instrumented ones unlikely and therefore instrument the whole bitcode file.
10 Related Work
@@
Much work has been put into side-channel analysis of programs for constant-time properties [2, 44, 14]. While there are still many open questions to be answered [20], we assume programs that are in line with the constant-time paradigm. There are several compilers that can verify constant-time properties during compilation [8, 32] as well as compilers that transform non constant-time code into constant-time code [12, 17]. Expanding an existing constant-time transforming compiler is possible in principle. In this work, we have opted for a standalone approach, to get precise measurements and reduce engineering effort. Under the assumption of being constant-time, further leakages like silent stores of secret-dependent memory writes [15] can be pinpointed.
Apart from analyzing programs for timing side-channel leakage, there are also leakage evaluations for examining leakage properties originating from upcoming microarchitectural optimizations [37, 7]. With the possibility of pinpointing leaking code pieces, many leakage mitigations approaches are proposed: On binary level, Cipherfix [42] is a ciphertext side-channel mitigation based on masking every memory write access with a randomly or pseudorandomly generated mask.
In the field of compiler level leakage defenses, cio [19] has been proposed as a mitigation against instruction-centric leakages, including silent store leakage. Regarding silent stores, cio rewrites each memory write so that there is an additional memory write with a different value before each possible memory write, so that no write can be suppressed because of equal values.
@@
Obelix [43] is a tool that combines side-channel leakage mitigations against many side-channels at once. Obelix also provides a protection level that mitigates ciphertext side-channel leakage but produces overhead factors around 1000×1000$\times$1000 × compared to the baseline execution and does not scale well to larger applications. In case only specific features are supported by the architecture, mitigations introducing lower overhead can be applied [46]; for open architectures even with instruction set architecture (ISA) augmentations [11, 45]. However, we assume a cloud setting where we cannot rely on ISA augmentations, making the proposed approaches inapplicable.
@@
By enabling hardware configurations that disable certain optimizations, certain side-channel leakages can be avoided. For operand-independent timing of instructions, ARM introduced DIT [6] and Intel DOIT [22]. The operand-independent timing modes disable data memory-dependent prefetching [21, 13]. Yet, the actual implementation is unclear and if the hypervisor can adjust the modes, software solutions are still needed as a backup.
@@
Protecting data in memory has not only been implemented with interleaving data on compiler level or masking data on binary level, but also with data space randomization (DSR) [10]. DSR uses randomization of data in memory via masks to hinder control-flow hijacking attacks. An extension of DSR is CoDaRR: The data does not only get randomized once but also re-randomized in tight intervals [31]. However, the overhead from such a re-randomization for all data does not scale for ciphertext side-channel protection with recurring memory writes. Schemes like Data and Pointer Prioritization (DPP) [1] that try optimizing similar overheads scale well only when their protection is restricted to data that might be subject to memory errors and associated unauthorized memory writes. Works like [29] and [30] promise selective encryption for data in memory. Yet, the encryption mode does not include freshness and thereby does not add protection for multiple memory writes with the same value to the same address as it would be needed for ciphertext side-channel mitigation; adding key updates introduces the common problem of randomness generation.
11 Conclusion
@@
In this work, we have studied memory-centric side-channel leakages that render constant-time code guarantees ineffective. We have discussed prerequisites for efficient ciphertext side-channel mitigations and implemented Zebrafix. We have shown that, while Zebrafix and interleaving in general can improve both the performance and security guarantees of ciphertext side-channel mitigations, this may come with practical constraints such as source code adjustments. Moreover, we have leveraged interleaving to devise mitigations against more generic memory-centric side-channel leakages like silent stores or data memory-dependent prefetching, and explained how holistic defenses can be developed. We conclude that, despite its limitations, interleaving is a promising approach to counter ciphertext side-channels and silent stores while solving major technical problems is needed to counter leakage from data memory-dependent prefetchers.
@@ Acknowledgment
Data Availability
Our source code is available on GitHub at the following URL: https://github.com/UzL-ITS/zebrafix. The repository includes a fork of LLVM with changes to the register allocator and interleaving support. Moreover, the repository contains the out-of-tree LLVM passes for the Zebrafix mitigation implementation, the memory tracer Pintool, and the libraries and benchmark code for the performance evaluation.

### v2

@@ VII-B Comparison of Silent Store Mitigations
The Zebrafix implementation protects against leakage from silent stores without requiring further adaptions as long as the silencing of writes cannot be decomposed into smaller chunks. We therefore compare Zebrafix to the silent store mitigation cio [15] that has also been implemented for x86-64. The Zebrafix implementation addresses the general root cause of memory-centric leakages. In line with that, the proof-of-concept implementation does not include a tracking for leaking or secret-dependent instructions, but rather instruments all memory writes. The silent store-only version of cio only covers a subset of all writes that are instrumented in Zebrafix. Zebrafix does, in contrast to cio, not have an automated verification mechanism but rather introduces manual verification of the non-leakiness of memory writes that are introduced by the register allocation.
@@
Out of the targets evaluated in cio, we compare against $\mathtt{Ed25519}$ and $\mathtt{ChaCha20}$ which were shown to be (partially) vulnerable in [7]. For $\mathtt{ChaCha20}$, we refer to the encryption; the decryption results are similar. The remaining targets evaluated in cio are not included in this case study because they contain instructions that are not supported in our proof-of-concept due to the way the interleaving is implemented based on vector register writes. Yet, the analyzed primitives are implemented to be executed in vector registers so that the leaking points are initialization and memory cleanup. It is possible to manually safeguard this limited set of code.
@@
The performance results for cio are based on results reported in their paper. The Zebrafix measurement setup equals the one described in Section VI-B. In line with cio, we use the $\mathtt{libsodium}$ reference implementations and protect stores of up to 8 byte of data. Protecting a binary against silent store leakage with cio introduces an overhead of $1.85$$\times$ for $\mathtt{ChaCha20}$ and an overhead of $3.76$$\times$ for $\mathtt{Ed25519}$. In contrast to that, the ciphertext side-channel protection from Zebrafix introduces an overhead factor of $2.1$$\times$ for $\mathtt{ChaCha20}$ and a factor of $1.6$$\times$ for $\mathtt{Ed25519}$.
@@
In line with [7], we found certain possible leakages in the $\mathtt{ChaCha20}$ target. However, those leakages are not originated in multiple systematic collisions as in $\mathtt{cmov}$ or $\mathtt{cswap}$ functionality but rather single collisions with a large set of possible values that are written during each execution. According to current knowledge, we do not deem those writes as exploitable which is also in line with cio being able to prune a large set of consecutive memory writes from their instrumentation.
@@
VIII Lessons Learned for Future Mitigations
@@
As described in Section IV, instrumenting on compiler middle end level has multiple advantages and disadvantages. One of the main disadvantages of any static instrumentation method is that language features, such as determining the size of a buffer by subtracting pointers without knowing the content type, cannot be instrumented semantically correctly, and must be manually replaced in the source code. Another common issue during our implementation was the evaluation of $\mathtt{sizeof}$, which already happens in the compiler front end. This occurred in cases where $\mathtt{sizeof}$ was used to determine the input size of operations like $\mathtt{memcpy}$ or $\mathtt{memset}$, which work on arrays. In our proof-of-concept implementation of Zebrafix, we work around that problem by retrieving the correct byte count from the LLVM $\mathtt{DataLayout}$ object and combining that with type information from the SVF pass.
We support memory writes up to 8 bytes in size, which is the integer size on x86-64. For larger writes (e.g., vector instructions), the data would have to be split into chunks and written within multiple interleaved chunks. Although we have not implemented vector support in Zebrafix, we do not expect any specific performance impact when rewriting vectorized stores. In contrast to the typical dynamic linking of cryptographic libraries, the interleaving-based leakage mitigation requires that either all dependencies are recompiled and statically linked or that the data from the instrumented code is properly handed over to unprotected functionality after declassification. The extensive changes to the memory layout cannot be handled transparently by non-instrumented code.
@@
Most cryptographic libraries are implemented in C, a language that lacks type safety, which poses certain key challenges for memory layout adjustments. Some manual intervention is needed for functions that use “indirect” casts on input data like putting a char array into a function that expects to work with 4-byte integers. Such casts might lead to semantically wrong results when applying operations that depend on the data type’s width like copying data from one buffer into another. The default handling case can only assume 8-byte data chunks in order to avoid out-of-bounds reads or writes, so any function that displays a warning has to be checked and then adjusted in the LLVM SVF pass handling for the correct pointer tracking. The currently implemented solution relies on the fact that function names are available in such cases for easing manual adjustments. With C, there is the choice between manual code adaptation, significant compiler code analysis, or extensive runtime type tracking. Alternatively, one can rely on a language in which such casts do not occur, and binary types of variables are always known.
@@
Another general issue arising in the proposed instrumentation approach stems from global variable usage: Either all code that uses global variables is instrumented or no code at all. In case a mitigation is implemented with additional secrecy tracking, the correct functionality cannot be guaranteed for global variable usage. If instrumented and non-instrumented parts of the code used global variables, the global variable state might destroy synchronization of values between different functions. Moreover, the general functionality can be compromised as even if all global variable usages were replaced with the instrumented version, there are usages depending on the global variables that cause problems in function parameters or functions like $\mathtt{memcpy}$ or $\mathtt{memset}$ or in $\mathtt{GEP}$ accesses for address calculations. [^5] Even with optimization level $\mathtt{-O2}$ and having all GEP accesses combined in one instruction due to the LLVM InstCombine pass, the iteration steps in arrays can still be wrong if the source element type of a GEP is not interleaved while the pointer of it points to an interleaved (instrumented) global variable. Further analyses would incur large overheads and make the resulting binary less stable; Apart from that, we deem the scenario of needing a non-instrumented version of each function alongside of instrumented ones unlikely and therefore instrument the whole bitcode file.
IX Related Work
@@
Much work has been put into analysis of programs for constant-time properties [2, 39, 11]. While there are still many open questions left [16], we assume programs that are in line with the constant-time paradigm. Under the assumption of being constant-time, further leakages like silent stores of secret-dependent memory writes [12] can be pinpointed.
@@
Apart from analyzing programs for timing side-channel leakage, there are also leakage evaluations for examining leakage properties originating from upcoming microarchitectural optimizations [32, 7]. With the possibility of pinpointing leaking code pieces, many leakage mitigations approaches are proposed: Obelix [38] is an ORAM-based tool that combines many side-channel leakage mitigations. Obelix also provides a protection level that mitigates ciphertext side-channel leakage through data rotation and interleaving but produces overhead factors around 1000×1000$\times$1000 × and does not scale to larger applications. In case only specific features are supported by the architecture, mitigations introducing lower overhead can be applied [41]; for open architectures even with instruction set architecture (ISA) augmentations [9, 40]. However, we assume a cloud setting where we cannot rely on ISA augmentations, making the proposed approaches inapplicable.
@@
By enabling hardware configurations that disable certain optimizations, certain side-channel leakages can be avoided. For operand-independent timing of instructions, ARM introduced DIT [6] and Intel DOIT [18]. The operand-independent timing modes disable data memory-dependent prefetching [17, 10]. Yet, the actual implementation is unclear and if the hypervisor can adjust the modes, software solutions are still needed as a backup. A discussion of potential software-based solutions is given in Appendix A.
@@
An extension of DSR is CoDaRR where the data gets re-randomized in tight intervals [27]. However, the overhead from such a re-randomization for all data does not scale for ciphertext side-channel protection with recurring memory writes. Schemes like Data and Pointer Prioritization (DPP) [1] that try optimizing similar overheads scale well only when their protection is restricted to data that might be subject to memory errors and associated unauthorized memory writes. Works like [25] and [26] promise selective encryption for data in memory. Yet, the encryption mode does not include freshness and thus does not add protection for multiple memory writes with the same value to the same address as needed for ciphertext side-channel mitigation; adding key updates introduces the common problem of randomness generation.
X Conclusion
@@
In this work, we have studied memory-centric side-channel leakages that render constant-time code guarantees ineffective. We have discussed prerequisites for efficient ciphertext side-channel mitigations and implemented Zebrafix. We have shown that Zebrafix and interleaving in general can improve both the performance and security guarantees of ciphertext side-channel mitigations, though this may come with practical constraints such as source code adjustments. Moreover, we have leveraged interleaving to devise mitigations against more generic memory-centric side-channel leakages like silent stores, and explained how holistic defenses can be developed. We conclude that, despite its limitations, interleaving is a promising approach to counter ciphertext side-channels and silent stores.
@@ Acknowledgment
Appendix A Data Memory-Dependent Prefetchers
Apart from the previously discussed leakages originating from ciphertext side-channels and silent stores, data memory-dependent prefetchers (DMPs) also fall into the category of memory-centric side-channels. In this section, we introduce the DMP leakage patterns and discuss to what extent interleaving can be used as a mitigation against the leakage.
A-A Leakage from Data Memory-Dependent Prefetchers
A-B Preventing Leakages from Data Memory-Dependent Prefetchers via Interleaving
With suitable instrumentation in place, all memory-centric leakages that arise from writing repetitive data to memory (see Section II) can be avoided. However, there also are tight relations between ciphertext side-channels and read-based leakages like data memory-dependent prefetching (DMP). In this section, we analyze interconnections between ciphertext side-channel and DMP attacks, and outline to what extent interleaving may be adapted for countering DMP leakage. Our threat model assumption for these considerations is the same as in Section VII.
Ciphertext Side-Channel Mitigation vs. DMP
The key idea to preventing DMP leakage is breaking the observable relation between data and the corresponding DMP activation. Ciphertext side-channel mitigations that add masking to secret plaintexts before writing to memory are thus effective, if the data is randomized. However, for providing a sufficient security level, existing masking approaches like Cipherfix incur high overheads due to the expensive random number generation [37]. The 16-byte block interleaving in Zebrafix that promises lower performance overhead cannot be used directly: Pointers are typically aligned to 8-byte boundaries, so they are unaffected by the additional counter in every second 8-byte chunk. If the DMP still triggers for the new memory layout, it can directly dereference these pointers, ignoring the counters altogether.
Extending Interleaving to DMP
If interleaving is to be used for DMP prevention, a smaller granularity is needed to break up DMP activation patterns, effectively reducing the block size from 16 bytes to 8 bytes. Instead of dividing data into 8-byte chunks and interleaving them with 8-byte counters, the data and counter chunk sizes need to be shrunk to implementation-dependent smaller chunks, such as 4 bytes each. By that, the supported data size is reduced to 4 bytes, which causes significant problems: 8-byte integers, which are mandated by the architecture and widely used in cryptographic libraries, need to be split into two 4-byte parts. Such split integers can no longer be read through a standard memory load, but need to be emulated. Additionally, the changes affect a primitive type that is assumed to exist throughout the compiler framework, so tweaks need to be applied at all layers. Due to the significant engineering effort, we leave the implementation of the exploration of this interleaving adaption for future work.
In addition to supporting another block size, the former freshness value has to be replaced by a combination of counter and DMP invalidation. For pure DMP mitigation, the counters do not need to be incremented and can be set to a fixed value that prevents DMP activation. If the mitigation is meant also protect against other classes of memory-centric side-channel leakage, the counters can be incremented as before while ensuring that no values occur that again trigger the DMP. An illustration of the new memory layout can be found in Figure 10. An implementation needs to be careful with endianness, as to ensure that the counter ends up transforming the necessary bytes of a potential pointer.
In this mitigation description, we assume that only canonical user-level pointer candidates, i.e., virtual addresses that may trigger prefetching, leak secret information via getting dereferenced. This is in line with current findings [34], that exclude leakage of non-canonical or kernel space addresses. With interleaving-based methods, mitigations may targets the root cause of the leakage, i.e., the DMP activation on a data chunk. Thereby, the aim is to also prevent subsequent exploitable leakages stemming from side-channels based on address translation or TLB lookups [34]. In addition, we assume the worst case leakage scenario, i.e., that the microarchitecture is already in a state that would immediately leak information when a pointer candidate is selected for prefetching. Note that we base the explanation of extensibility on the attack vectors presented in [34], combined with theoretical reasoning about address translation properties. Concrete implementations of microarchitectural optimizations may vary regarding specific processor or architectural features. We therefore conclude that for an efficient and secure mitigation against DMP leakage, the usage of verifiable DOIT/DIT modes is more effective if those are available.